OpenStudy (anonymous):

A six-sided fair die is rolled 4 times in a row. The probability of getting a 4 only on the last trial is ?

OpenStudy (anonymous):

my guess was 4/6 but i got it wrong

OpenStudy (afrodiddle):

Chica, why would it be 4/6? o.o

OpenStudy (anonymous):

i dont know, i guessed

OpenStudy (welshfella):

you got to work out the scenario its not 4, not 4 , not 4, 4

OpenStudy (afrodiddle):

you were on the right track but I'm afraid you got a brain fart.

OpenStudy (welshfella):

P(4) = 1/6 , P(not 4) = 5/6 multiply these probabilities

OpenStudy (welshfella):

the throws are independent of each other so the probs are multiplied required probability = 5/6 * 5/6 * 5/6 * 1/6

OpenStudy (anonymous):

0.96....

OpenStudy (welshfella):

125/1296 or 0.096

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

thanks